# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ Abstract.
First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD) have been proven effective in training neural networks. In the setting of over-parameterization, there is a line of work demonstrating that randomly initialized (stochastic) gradient descent converges to a globally optimal solution at a linear convergence rate for the quadratic loss function. However, the learning rate of GD in training two-layer neural networks has a poor dependence on the sample size and the Gram matrix, resulting in a slow training process. In this paper, we show that for the $L^{2}$ regression problems, the learning rate can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$ to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, which implies that GD enjoys a faster convergence rate. Moreover, we further generalize the method for GD in training two-layer Physics-Informed Neural Networks (PINNs), showing a similar improvement for the learning rate. Although the improved learning rate depends mildly on the Gram matrix, we still need to set it small enough in practice due to the agnostic eigenvalues of the Gram matrix. More importantly, the convergence rate relies on the least eigenvalue of the Gram matrix, leading to slow convergence. In this work, we provide the convergence analysis of natural gradient descent (NGD) in training two-layer PINNs. We show that the learning rate can be $\mathcal{O}(1)$ and at this time, the convergence rate is independent of the Gram matrix.
@@ 1. Introduction
In recent years, neural networks have achieved remarkable breakthroughs in the fields of image recognition [1], natural language procssing [2], reinforcement learning [3] and so on. Moreover, due to the flexibility and scalability of neural networks, researchers are paying much attention in exloring new methods involving neural networks for handling problems in scientific computing, where one long-standing and essential problem is solving partial differntial equations (PDEs) numerically. Classical numerical methods, such as finite difference, finite volume and finite elements methods, suffer the curse of dimensionality in solving high-dimensional PDEs. Due to this drawback, various methods involving neural networks have been proposed for solving different type PDEs [4, 5, 6, 7, 8]. Among them, the most representative approach is Physics-Informed Neural Networks (PINNs) [5]. In the framework of PINNs, one incorporate PDE constraints into the loss function and train the neural network with it. With the use of automatic differentiation, the neural network can be efficiently trained by first-order or second-order methods.
@@
In the applications of neural networks, one inevitable issue is the selection of the optimization method. First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD), are widely used in optimizing neural networks as they only calculate the gradient and therefore are computationally efficient. Beside the first-order methods, there has been much interest in utilizing second-order optimization methods to speed up the training, not only for the regression problems [9] but also for the problems related to PDEs [4, 5].
@@
For the convergence part of the optimization method, it has been shown that gradient descent algorithm can even achieve zero training loss under the setting of over-parametrization, which refers to a situation where a model has more parameters than necessary to fit the data [10, 11, 12, 13, 14, 15]. These works are based on the idea of neural tangent kernel (NTK), which shows that training multi-layer fully-connected neural networks via gradient descent is equivalent to performing a certain kernel method as the width of every layer goes to infinity. As for the finite width neural networks, with more refined analysis, it can be shown that the parameters are closed to the initialization throughout the entire training process when the width is large enough. It directly leads to the linear convergence for GD. Despite these beautiful convergence results, the learning rate depends on the sample size and the Gram matrix. Thus, it need to be sufficiently small to guarantee the convergence in practice. But doing so results in a slow training process. In contrast to the first-order methods, the second-order methods enjoy fast convergence due to the absence of requirements for the learning rate [16].
@@ 1.1. Contributions
  • For the $L^{2}$ regression problems, we demonstrate that the learning rate $\eta$ of gradient descent can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$, as shown in [10], to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the ReLU activation function and the random initialization, $\lambda_{0}$ is the least eigenvalue of $\bm{H}^{\infty}$. [17] has also shown the same improvement of the learning rate, but it requires that $m=\Omega(\frac{n^{6}}{\lambda_{0}^{4}\delta^{3}})$. Moreover in [18], the dependence on $n$, i.e. $\Omega(n^{6})$, is necessary and can not be improved due to the requirement of the proof method. Different from the method in [17], our method comes from a new recursion formula of gradient descent and it can be easily generalized to PINNs. Compared to [17], we only require that $m=\Omega(\frac{n^{4}}{\lambda_{0}^{4}}(\log(\frac{n}{\delta}))^{2})$.
  • For the PINNs, we simultaneously improve both the learning rate $\eta$ of gradient descent and the requirement of the width $m$. The improvements rely on a new recursion formula of gradient descent, which is similar to that for regression problems. Specifically, we can improve the learning rate $\eta=\mathcal{O}(\lambda_{0})$ required in [19] to $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$ and the requirement for the width $m$, i.e. $m=\tilde{\Omega}(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}})$, can be improved to $m=\tilde{\Omega}(\frac{1}{\lambda_{0}^{4}}(\log(\frac{n_{1}+n_{2}}{\delta})))$.
  • We provide the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. Compared to gradient descent, the convergence rate of NGD is independent of $\lambda_{0}$, which results in a faster convergence. Meanwhile, we show that the learning rate can be even $\mathcal{O}(1)$. When the activation function is smooth, NGD enjoys quadratic convergence rate.
@@ 1.2. Related Works
First-order methods. There are mainly two ways to study the optimization of neural networks and try to understand why first-order methods can find a global minimum. One way is to analyse the optimization landscape [15, 20]. It has been shown that gradient descent can find a global minimum in polynomial time if the optimization landscape has some good geometric properties. However, some unrealistic assumptions in these works make it difficult to generalize to practical neural networks. Another way to understand the optimization of neural networks is to analyze the optimization dynamics of first-order methods. For the two-layer ReLU neural networks, as shown in [10], as long as the width $m$ is large enough and no two inputs are parallel, randomly initialized gradient descent converges to a globally optimal solution at a linear rate. Later, the results have been generalized to deep neural networks with smooth activation functions [11]. Both rely on that the Gram matrices are stable throughout the training process, which leads to the convergence to the global minimum. Besides regression and classification problems, [19] demonstrated the convergence of the gradient descent for two-layer PINNs by the similar analysis of optimization dynamics. However, both [10] and [19] require that the learning rate be sufficiently small to converge. In this work, we conduct a refined analysis of gradient descent for $L^{2}$ regression problems and PINNs, leading to a milder requirement for the learning rate.
Second-order methods. Although second-order methods possess better convergence rate, they are rarely used in training deep neural networks due to the prohibitive computational cost. As a variant of the original Gauss-Newton method, natural gradient descent (NGD) is more efficient in practice. Meanwhile, as shown in [21] and [16], it also enjoys faster convergence rate for the $L^{2}$ regression problems than of gradient descent. In this paper, we provide the convergence analysis for NGD in training two-layer PINNs.
@@ 1.3. Notations
We denote $[n]=\{1,2,\cdots,n\}$ for $n\in\mathbb{N}$. Given a set $S$, we denote the uniform distribution on $S$ by $Unif\{S\}$. We use $I\{E\}$ to denote the indicator function of the event $E$. For variables $a$ and $b$, $a\lesssim b$ is equivalent to $a\leq Cb$ for some universal constant $C$. For two functions $f_{1}(n)$ and $f_{2}(n)$, the relation $f_{1}(n)=\mathcal{O}(f_{2}(n))$ is equivalent to $f_{1}(n)\leq Cf_{2}(n)$ for some universal constant $C$. And similarly, $f_{1}(n)=\Omega(f_{2}(n))$ is equivalent to $f_{1}(n)\geq Cf_{2}(n)$. Throughout the paper, we use boldface denote vectors. Given $x_{1},\cdots,x_{d}\in\mathbb{R}$, we use $(x_{1},\cdots,x_{d})$ or $[x_{1},\cdots,x_{d}]$ to denote a row vector with $i$-th component $x_{i}$ for $i\in[d]$ and then $(x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d}$ is a column vector.
@@ 1.4. Organization of this Paper
We present the improvements of the learning rate of gradient descent for $L^{2}$ regression problems and PINNs in Section 2 and Section 3 respectively. In Section 4, we show the convergence of natural gradient descent for training PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. We conclude in Section 5 and the detailed proofs are put in the Appendix for readability and brevity.
2. Improved Learning Rate of Gradient Descent for $L^{2}$ Regression Problems
2.1. Problem Setup
In this section, we consider a two-layer neural network $f$ with the following form.
$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\tilde{\bm{x}}),$ (1)
where $\bm{x}\in\mathbb{R}^{d}$ is the input, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$, $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T},\bm{a}=(a_{1},\cdots,a_{m})^{T}$, for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+1}$ is the weight vector of the first layer, $a_{r}\in\mathbb{R}$ is the output weight and $\sigma(\cdot)$ is the ReLU activation function. Different from the setup in [10], the two-layer neural network that we consider has bias term. In the following, we assume that each input vector $\bm{x}\in\mathbb{R}^{d}$ has been augmented to $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$. With a little abuse of notation, we write $\bm{x}$ for $\tilde{\bm{x}}$ and write $f$ as
$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}).$
For the $L^{2}$ regression problem, given training data set $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, the aim is to minimize the loss function
$L(\bm{w},\bm{a}):=\sum\limits_{i=1}^{n}\frac{1}{2}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})^{2}.$ (2)
Before training, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$ and output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$. In the training process, we fix the second layer and optimize the first layer by gradient descent (GD), i.e., we update the weights by the following formulation.
$\bm{w}(k+1)=\bm{w}(k)-\eta\frac{\partial L(\bm{w}(k))}{\partial\bm{w}},$ (3)
where $\eta>0$ is the learning rate, $k\in\mathbb{N}$ and $L(\bm{w}(k))$ is an abbreviation of $L(\bm{w}(k),\bm{a})$. From the form of $L(\bm{w})$, we can deduce that
$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{m}}\sum\limits_{i=1}^{n}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})\bm{x}_{i}I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}.$ (4)
We denote $u_{i}(k)=f(\bm{x}_{i};\bm{w}(k),\bm{a})$ the prediction on input $\bm{x}_{i}$ at $k$-th iteration and we let $\bm{u}(k)=(u_{1}(k),\cdots,u_{n}(k))^{T}\in\mathbb{R}^{n}$ be the prediction vector at $k$-th iteration. For the labels, we define $\bm{y}:=(y_{1},\cdots,y_{n})^{T}\in\mathbb{R}^{n}$.
According to [10], in the continuous setting, the dynamics of predictions can be written as
$\frac{d}{dt}\bm{u}(t)=\bm{H}(t)(\bm{y}-\bm{u}(t)),$ (5)
where $H(t)\in\mathbb{R}^{n\times n}$ is the Gram matrix at time $t$ with $(i,j)$-entry
$H_{ij}(t)=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}(t)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(t)^{T}\bm{x}_{j}\geq 0\right\}.$ (6)
In the setting of over-parameterization and randomly initialization, [10] has shown that (1) $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(\sqrt{1/m})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the initialization with $(i,j)$-th entry
$H_{ij}^{\infty}=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}^{T}\bm{x}_{i}\geq 0,\bm{w}^{T}\bm{x}_{j}\geq 0\right\}\right]$ (7)
and (2) $\|\bm{H}(t)-\bm{H}(0)\|_{2}=\mathcal{O}(\sqrt{1/m})$ for all $t>0$. Therefore, as $m\to\infty$, the dynamics of the predictions are charactered by $H^{\infty}$, leading to the linear convergence.
2.2. Main Results
Assumption 1.
For $i\in[n]$, $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$ and $|y_{i}|\leq 1$, where $\bm{x}_{i}\in\mathbb{R}^{d+1}$ is the augmented input.
Assumption 2.
No two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are parallel, i.e., for any $\bm{x}_{t},\bm{x}_{s}\in\{\bm{x}_{i}\}_{i=1}^{n}$ and any $\alpha\in\mathbb{R}$, we have $\bm{x}_{t}\neq\bm{x}_{s}$.
Since inputs are all augmented, Assumption 2 is equivalent to that no two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are equal, which holds naturally. Under Assumption 2, Theorem 3.1 in [10] implies that $\lambda_{0}:=\lambda_{min}(\bm{H}^{\infty})>0$, which is crucial in the convergence analysis.
As stated before, there are two improtant facts that determine the optimization dynamics, one is that at initialization $\bm{H}(0)$ is closed to $\bm{H}^{\infty}$ and another is that $\bm{H}(k)$ does not go far away from the initialization $\bm{H}(0)$ for all $k\in\mathbb{N}$. These facts are supported by the following two lemmas.
Lemma 1 (Lemma 3.1 in [10]).
If $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$, we have with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3\lambda_{0}}{4}$.
Lemma 2.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-n^{2}e^{-mR/4}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then the matrix $\bm{H}(\bm{w})\in\mathbb{R}^{n\times n}$ defined by
$H(\bm{w})_{ij}=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}.$
satisfies
$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}<4nR.$ (8)
Remark 1.
Although Lemma 2 looks almost same to Lemma 3.1 in [18], the proof of Lemma 3.1 in [18] lacks an important part, i.e., the proof of the measurability for the related variable. In fact, Lemma 2 ensures us to bound the change of $\bm{H}(\bm{w})$ when $\bm{w}$ in a small ball. However, the set consisting of vectors in the small ball is uncountable, thus the related variable may not be measurable due to the discontinuity of the indicator function appeared in the definition of $\bm{H}(\bm{w})$. In the proof, we borrow the concept of pointwise measurability in the empirical process (see Chapter 8 in [22]) to bridge the gap. Then we can bound the supremum with a new random variable. By applying a concentration inequality to this new random variable, we can reach the conclusion.
Abount the convergence of gradient descent, [10] has demonstrated that if the learning rate $\eta=\mathcal{O}(\lambda_{0}/n^{2})$, then randomly initialized gradient descent converges to a globally optimal solution at a linear convergence rate when $m$ is large enough. The requirement of $\eta$ is from the decomposition for the residual in the $(k+1)$-th iteration, i.e.,
$\bm{y}-\bm{u}(k+1)=\bm{y}-\bm{u}(k)-(\bm{u}(k+1)-\bm{u}(k)).$ (9)
Instead of decomposing the residual into the two terms as above, we write it as follows, which serves as a recursion formula.
Lemma 3.
For all $k\in\mathbb{N}$, we have
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta H(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k),$ (10)
where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n}(k))^{T}\in\mathbb{R}^{n}$ with $i$-th
$I_{1}^{i}(k):=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$ (11)
After [10], [17] improved the network size $m$ to $m=\mathcal{O}(\lambda_{0}^{-4}n^{4}\log^{3}(n/\delta)\log(m/\delta))$, but it still requires $\eta=\mathcal{O}(\lambda_{0}/n^{2})$. Another work [17] provided a refined analysis over [10], which shows that $\eta=\frac{1}{C_{1}\|\bm{H}^{\infty}\|_{2}}$ is enough for the convergence rate $1-\frac{\lambda_{0}C_{2}}{C_{1}\|\bm{H}{\infty}\|_{2}}$, but the constants $C_{1},C_{2}$ may depend on the parameters $\lambda_{0},n$ and $\delta$. It requires the network size $m$ to be $\Omega(\lambda_{0}^{-4}n^{6}\delta^{-3})$ to make $C_{1},C_{2}$ independent of $\lambda_{0},n$ and $\delta$. In this paper, we improve both the learning rate and network size and the main reslut is the following theorem.
Theorem 1.
Under Assumption 1 and Assumption 2, if we set the number of hidden nodes $m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right)$ and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
$\|\bm{y}-\bm{u}(k)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ (12)
for all $k\in\mathbb{N}$.
Remark 2.
Our proof method requires that $\bm{I}-\eta\bm{H}(k)$ is positive definite for all $k\in\mathbb{N}$. As $\bm{H}(k)$ is closed to $\bm{H}^{\infty}$, the requirement is satisfied if $\eta\|\bm{H}^{\infty}\|_{2}\lesssim 1$. Note that $\|\bm{H}^{\infty}\|_{2}\leq n$, it is sufficient to set $\eta=\mathcal{O}(1/n)$, which results in an improvement of $\mathcal{O}(\lambda_{0}/n)$.
3. Improved Learning Rate of Gradient Descent for Two-Layer Physics-Informed Neural Networks
3.1. Problem Setup

### v3

@@ Abstract.
In the context of over-parameterization, there is a line of work demonstrating that randomly initialized (stochastic) gradient descent (GD) converges to a globally optimal solution at a linear convergence rate for the quadratic loss function. However, the learning rate of GD for training two-layer neural networks exhibits poor dependence on the sample size and the Gram matrix, leading to a slow training process. In this paper, we show that for training two-layer $\text{ReLU}^{3}$ Physics-Informed Neural Networks (PINNs), the learning rate can be improved from $\mathcal{O}(\lambda_{0})$ to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, implying that GD actually enjoys a faster convergence rate. Despite such improvements, the convergence rate is still tied to the least eigenvalue of the Gram matrix, leading to slow convergence. We then develop the positive definiteness of Gram matrices with general smooth activation functions and provide the convergence analysis of natural gradient descent (NGD) in training two-layer PINNs, demonstrating that the learning rate can be $\mathcal{O}(1)$ and at this rate, the convergence rate is independent of the Gram matrix. In particular, for smooth activation functions, the convergence rate of NGD is quadratic. Numerical experiments are conducted to verify our theoretical results.
@@ 1. Introduction
In recent years, neural networks have achieved remarkable breakthroughs in the fields of image recognition [1], natural language processing [2], reinforcement learning [3], and so on. Moreover, due to the flexibility and scalability of neural networks, researchers are paying much attention in exploring new methods involving neural networks for handling problems in scientific computing. One long-standing and essential problem in this area is solving partial differential equations (PDEs) numerically. Classical numerical methods, such as finite difference, finite volume and finite elements methods, suffer from the curse of dimensionality when solving high-dimensional PDEs. Due to this drawback, various methods involving neural networks have been proposed for solving different type PDEs [4, 5, 6, 7, 8]. Among them, the most representative approach is Physics-Informed Neural Networks (PINNs) [5]. In the framework of PINNs, one incorporate PDE constraints into the loss function and train the neural network with it. With the use of automatic differentiation, the neural network can be efficiently trained by first-order or second-order methods.
@@
In the applications of neural networks, one inevitable issue is the selection of the optimization methods. First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD), are widely used in optimizing neural networks as they only calculate the gradient, making them computationally efficient. In addition to first-order methods, there has been significant interest in utilizing second-order optimization methods to accelerate training. These methods have proven to be applicable not only to regression problems, as demonstrated in [9], but also to problems related to PDEs, as shown in [4, 5].
@@
As for the convergence aspect of the optimization methods, it has been shown that gradient descent algorithm can even achieve zero training loss under the setting of over-parameterization, which refers to a situation where a model has more parameters than necessary to fit the data [10, 11, 12, 13, 14, 15, 16, 17]. These works are based on the idea of neural tangent kernel (NTK)[18], which shows that training multi-layer fully-connected neural networks via gradient descent is equivalent to performing a certain kernel method as the width of every layer goes to infinity. As for the finite width neural networks, with more refined analysis, it can be shown that the parameters are closed to the initializations throughout the entire training process when the width is large enough. This directly leads to the linear convergence for GD. Despite these attractive convergence results, the learning rate depends on the sample size and the Gram matrix, so it needs to be sufficiently small to guarantee convergence in practice. However, doing so results in a slow training process. In contrast to first-order methods, the second-order method natural gradient descent (NGD) has been shown to enjoy fast convergence for the $L^{2}$ regression problems, as demonstrated in [19, 20]. However, the convergence of NGD in the context of training PINNs is still an open problem. In this paper, we demonstrate that when training PINNs, NGD indeed enjoys a faster convergence rate.
@@ 1.1. Contributions
  • For the PINNs, we simultaneously improve both the learning rate $\eta$ of gradient descent and the requirement for the width $m$. The improvements rely on a new recursion formula for gradient descent, which is similar to that for regression problems. Specifically, we can improve the learning rate $\eta=\mathcal{O}(\lambda_{0})$ required in [21] to $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$ and the requirement for the width $m$, i.e. $m=\widetilde{\Omega}\left(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}}\right)$, can be improved to $m=\widetilde{\Omega}\left(\frac{1}{\lambda_{0}^{4}}(\log(\frac{n_{1}+n_{2}}{\delta}))\right)$, where $\widetilde{\Omega}$ indicates that some terms involving $\log(m)$ are omitted.
  • We present a framework for demonstrating the positive definiteness of Gram matrices for a variety of commonly used smooth activation functions, including the logistic function, softplus function, hyperbolic tangent function, and others. This conclusion is not only applicable to the PDE we have considered but can also be naturally extended to other forms of PDEs.
  • We provide the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. Due to the distinct optimization dynamics of NGD compared to GD, the learning rate can be $\mathcal{O}(1)$. Consequently, the convergence rate is independent of $n$ and $\lambda_{0}$, leading to faster convergence. Moreover, when the activation function is smooth, NGD can achieve a quadratic convergence rate.
@@ 1.2. Related Works
First-order optimizers. There are mainly two approaches to studying the optimization of neural networks and understanding why first-order methods can find a global minimum. One approach is to analyze the optimization landscape, as demonstrated in [22, 23]. It has been shown that gradient descent can find a global minimum in polynomial time if the optimization landscape possesses certain favorable geometric properties. However, some unrealistic assumptions in these works make it challenging to generalize the findings to practical neural networks. Another approach to understand the optimization of neural networks is by analyzing the optimization dynamics of first-order methods. For the two-layer ReLU neural networks, as shown in [10], randomly initialized gradient descent converges to a globally optimal solution at a linear rate, provided that the width $m$ is sufficiently large and no two inputs are parallel. Later, these results were extended to deep fully-connected feedforward neural networks and ResNet with smooth activation functions [11]. Results for both shallow and deep neural networks depend on the stability of the Gram matrices throughout the training process, which is crucial for convergence to the global minimum. In addition to regression and classification problems, [21] demonstrated the convergence of the gradient descent for two-layer PINNs through a similar analysis of optimization dynamics. However, both [10] and [21] require a sufficiently small learning rate and a large enough network width to achieve convergence. In this work, we conduct a refined analysis of gradient descent for PINNs, resulting in milder requirements for the learning rate and network width.
[…]
@@ 1.3. Notations
We denote $[n]=\{1,2,\cdots,n\}$ for $n\in\mathbb{N}$. Given a set $S$, we denote the uniform distribution on $S$ by $Unif\{S\}$. We use $I\{E\}$ to denote the indicator function of the event $E$. For two positive functions $f_{1}(n)$ and $f_{2}(n)$, we use $f_{1}(n)=\mathcal{O}(f_{2}(n))$, $f_{2}(n)=\Omega(f_{1}(n))$ or $f_{1}(n)\lesssim f_{2}(n)$ to represent $f_{1}(n)\leq Cf_{2}(n)$, where $C$ is a universal constant $C$. A universal constant means a constant independent of any variables. Throughout the paper, we use boldface to denote vectors. Given $x_{1},\cdots,x_{d}\in\mathbb{R}$, we use $(x_{1},\cdots,x_{d})$ or $[x_{1},\cdots,x_{d}]$ to denote a row vector with $i$-th component $x_{i}$ for $i\in[d]$ and then $(x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d}$ is a column vector.
@@ 1.4. Organization of this Paper
In Section 2, we provide the problem setup for training two-layer PINNs. We then present the improved convergence results of gradient descent for PINNs in Section 3. In Section 4, we analyze the convergence of natural gradient descent in training two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. In Section 5, we conduct experiments to verify the theoretical results. The limitations are briefly discussed in Section 6 and we conclude in Section 7. All the detailed proofs and experiments are provided in the Appendix for readability and brevity.
2. Problem Setup